# Reporting Risks in AI-based Assistive Technology Research: A Systematic Review

### v1

@@ 4.2.1 Question 1: Does the paper present a working demo or examples of how the system behaves?
As shown in figure 4, 85 papers have answered positively Question 1, meaning they have a working demo or an output of the system’s behaviour. This could be in the form of a picture of the system’s output or showing the specific results of a test case. The other 16 papers did not address this question, whether they do not represent any output or the question does not apply to them, for example, in cases of studies that are in the early steps or are proposals.
Figure 4: The number of papers in the combined set according to their answer to Question 1: 84 Yes, and 16 No
@@ 4.2.2 Question 2: Is there an evaluation of the system by conducting a human study?
Figure 5: The number of papers in the combined set according to their answer to Question 2: 28 Yes, and 72 No
@@ 4.2.3 Question 3: Is the human study ecologically valid for persons with visual impairment?
As shown in figure 6, among the papers that answered yes to Question 1, only 16/85 = 19% of them have had at least one participant with visual impairment for the experiments. This suggests that, in total, out of all the papers that were checked, less than half of them have been evaluated by persons from the sight-loss community who are the target users. So, the theoretical evaluation of the system does not represent the community targeted by the research. Also, compared to the answers to Question 2, five papers had human study, but none of the participants were from the sight-loss community, and sighted or blindfolded participants did the test. For example, Akkapusit and Ko [14] have conducted a human study with 20 participants of different ages, genders, and experience levels using the provided equipment for the evaluation. However, all participants were sighted and blindfolded for the study. This is not an appropriate approach to evaluate an assistive technology intended to be used by the sight-loss community.
Figure 6: The number of papers in the combined set according to their answer to Question 3: 23 Yes, and 77 No
@@ 4.2.4 Question 4: Does the paper consider any threats to the validity of the human study?
Figure 8: The number of papers in the combined set according to their answer to Question 5: 78 Yes, and 22 No
@@ 4.2.6 Question 6: Does the paper report examples of failures?
Figure 9: The number of papers in the combined set according to their answer to Question 6: 25 Yes, and 75 No
@@ 4.2.7 Question 7: Does the paper give specific information about when and how the system will fail?
Figure 10: The number of papers in the combined set according to their answer to Question 7: 34 Yes, and 66 No
@@ 4.2.8 Question 8: Does the paper talk about the consequences of the reported/systems failures?
Figure 11: The number of papers in the combined set according to their answer to Question 8: 3 Yes, and 97 No
Discussion
Most studies demonstrated a working example of their system’s behaviour; however, among these papers, only a few evaluated their working on an actual test with human participants. Even among those who have a human experiment, a few papers did not have participants with visual impairment. This and many other factors mentioned in 4.2.4 question the validity of human studies, but very few papers talk about threats to the validity of their human study.
On the failure consideration side, most studies acknowledged that their system could create incorrect outputs (Question 5). However, again, only a tiny percentage of these papers specified the situations in which the system might fail or gave examples of the system’s incorrect results. Moreover, By analyzing the review results, we noticed that among the papers without human studies, only 13.7% of them reported the system’s failure cases. In comparison, among the papers with human studies, 44.4% reported examples of their system’s failure. This number is 50% for the studies that had a human study with participants from the sight-loss community. This comparison suggests that without the presence of a human study, the chance of having unnoticed failure cases increases.
In the end, only a few papers give awareness about what might happen due to the system’s failure to clear the associated risks of using the product.
@@ Discussion
Limitations
A limitation of this work is narrowing down our context to academic studies and published papers. We acknowledge that there are assistive technologies invented or provided by companies that do not publish in academic venues. Still, as stated in our research question, we were primarily interested here in exploring academia and research on AI-based assistive technologies.
Furthermore, the final search string was arrived at in order to be comprehensive while avoiding excessive irrelevant results. Still, no perfect search string could cover all the related papers. So, it is possible that some papers are not captured with our search string and are not included in our search results.
@@ 5 Conclusion
In this work, we conducted a systematic literature analysis of the published studies to understand how the failures and risks of AI-based assistive technologies for persons with visual impairment are presented in academia. We framed eight research questions, which were then applied to a subset of 100 papers randomly sampled from the total number of 648 papers in the systematic search results. The questions captured both human study and risk analysis aspects of the study. Results show only a few papers (19%) tested their work in a human study with visually impaired participants. Moreover, there is less information on the cases of failures and possible risks of introduced systems. We validated the method as the search string and inclusion/exclusion criteria effectively captured a considerable amount of relevant literature within our research scope, and the proposed analysis questions were answerable and provided meaningful results.
In our future work, we look to take the findings of this research to introduce guidelines for researchers on how to evaluate AI-based assistive technology well.
This will ultimately empower people to judge the trustworthiness of a newly presented assistive technology.

### v2

@@ 4.2.1 Question 1: Does the paper present a working demo or examples of how the system behaves?
As shown in figure 4, 85 papers have answered positively Question 1, meaning they have a working demo or an output of the system’s behaviour. This could be in the form of a picture of the system’s output or showing the specific results of a test case. The other 15 papers did not address this question, whether they do not represent any output or the question does not apply to them, for example, in cases of studies that are in the early steps or are proposals.
Figure 4: The number of papers in the combined set according to their answer to Question 1: 85 Yes, and 15 No
@@ 4.2.2 Question 2: Is there an evaluation of the system by conducting a human study?
Figure 5: The number of papers in the combined set according to their answer to Question 2: 27 Yes, and 73 No
@@ 4.2.3 Question 3: Is the human study ecologically valid for persons with visual impairment?
As shown in figure 6, among the papers that answered yes to Question 1, only 16/85 = 19% of them have had at least one participant with visual impairment for the experiments. This suggests that, in total, out of all the papers that were checked, only a few of them have been evaluated by persons from the sight-loss community who are the target users. So, the theoretical evaluation of the system does not represent the community targeted by the research. Also, compared to the answers to Question 2, five papers had human study, but none of the participants were from the sight-loss community, and sighted or blindfolded participants did the test. For example, Akkapusit and Ko [14] have conducted a human study with 20 participants of different ages, genders, and experience levels using the provided equipment for the evaluation. However, all participants were sighted and blindfolded for the study. This is not an appropriate approach to evaluate an assistive technology intended to be used by the sight-loss community.
Figure 6: The number of papers in the combined set according to their answer to Question 3: 16 Yes, and 84 No
@@ 4.2.4 Question 4: Does the paper consider any threats to the validity of the human study?
Figure 8: The number of papers in the combined set according to their answer to Question 5: 74 Yes, and 26 No
@@ 4.2.6 Question 6: Does the paper report examples of failures?
Figure 9: The number of papers in the combined set according to their answer to Question 6: 22 Yes, and 78 No
@@ 4.2.7 Question 7: Does the paper give specific information about when and how the system will fail?
Figure 10: The number of papers in the combined set according to their answer to Question 7: 32 Yes, and 68 No
@@ 4.2.8 Question 8: Does the paper talk about the consequences of the reported/systems failures?
Figure 11: The number of papers in the combined set according to their answer to Question 8: 2 Yes, and 98 No
5 Discussion
Most studies demonstrated a working example of their system’s behaviour; however, among these papers, only a few evaluated their working on an actual test with human participants. Even among those who have a human experiment, some papers had blindfolded participants instead of persons with visual impairment. This and many other factors mentioned in 4.2.4 question the validity of human studies, but very few papers talk about threats to the validity of their human study.
On the failure consideration side, most studies acknowledged that their system could create incorrect outputs (Question 5). However, again, only a tiny percentage of these papers specified the situations in which the system might fail or gave examples of the system’s incorrect results. Moreover, By analyzing the review results, we noticed that among the papers without human studies, only 13.7% of them reported the system’s failure cases. In comparison, among the papers with human studies, 44.4% reported examples of their system’s failure. This number is 50% for the studies that had a human study with participants from the sight-loss community. This comparison suggests that without the presence of a human study, the chance of having unnoticed failure cases increases. In the end, only a few papers give awareness about what might happen due to the system’s failure to clear the associated risks of using the product.
@@ 5 Discussion
6 Limitations
A limitation of this work is narrowing down our context to academic studies and published papers. We acknowledge that there are assistive technologies invented or provided by companies that do not publish in academic venues. Still, as stated in our research question, we were primarily interested here in exploring academia and research on AI-based assistive technologies. Furthermore, the final search string was arrived at in order to be comprehensive while avoiding excessive irrelevant results. Still, no perfect search string could cover all the related papers. So, it is possible that some papers are not captured with our search string and are not included in our search results.
@@ 7 Conclusion
In this work, we conducted a systematic literature analysis of the published studies to understand how the failures and risks of AI-based assistive technologies for persons with visual impairment are presented in academia. We framed eight research questions, which were then applied to a subset of 100 papers randomly sampled from the total number of 648 papers in the systematic search results. The questions captured both human study and risk analysis aspects of the study. Results show only a few papers (19% among those that have a prototype or demo of the system) tested their work in a human study with visually impaired participants. Moreover, there is less information on the cases of failures and possible risks of introduced systems. We validated the method as the search string and inclusion/exclusion criteria effectively captured a considerable amount of relevant literature within our research scope, and the proposed analysis questions were answerable and provided meaningful results.
In our future work, we look to take the findings of this research to introduce guidelines for researchers on how to evaluate AI-based assistive technology well. This will ultimately empower people to judge the trustworthiness of a newly presented assistive technology.